please help me solve this problem 2x to the second power -3=1
Answer by jillfired(3)

Since -3 does not contain the variable
variable to solve for, move it to the right-hand side of the equation
equation
2x^2=3=1
A mathematical statement that says two expressions have the same value; any number sentence with an = . by adding 3 to both sides.
Add 1 to 3 to get 4.
2x^2=4

Divide each term in the equation by 2
2x^2/2 = 4/2
Simplify the left side by cancelling out common factors:
A mathematical statement that says two expressions have the same value; any number sentence with an = . by 2.
x^2=4/2
simplify the right hand side
x^2=2
eliminate exponent by taking sqrt of both sides
so x= sqrt+/-2